# Joint Modelling of Astrophysical Systematics for Cosmology with LSST

### v2

@@ 4.3 Impact of varying luminosity function parameters in IA modelling only
Before considering a full joint modelling of IA and source photometric redshift distributions via the luminosity function, we consider also the impact of simply marginalising over the luminosity function parameters of the IA modelling an LSST cosmic shear analysis. In the case of the LSST Science Requirements Document (SRD) analysis of Mandelbaum et al. (2018) (which we use extensively in this work as a comparison point and an example of a standard analysis set-up), a Schechter luminosity function with fixed parameter values from the GAMA and DEEP2 surveys is used. The idea of including luminosity function parameters as nuisance IA modelling parameters was explored in Krause et al. (2015) for an LSST-like survey; here we present the results of doing so in our particular set-up to build intuition prior to introducing the luminosity function jointly within the redshift distribution and IA modelling.
@@
Since we aim to underscore the importance of including luminosity function parameters in the forecast, we analyze the resulting Fisher matrices to extract the $1-\sigma$ errors on parameters. To quantitatively assess the impact of this inclusion, we compute the fractional difference between the SRD and SRD+LF scenarios for these errors on the parameters. This is defined as:
@@
In Eq. 31, $\sigma$ represents a $1\sigma$ error on parameter $\theta$, where $X$ and $Y$ denote the forecasting scenarios, with $Y$ being the benchmark. This approach allows us to specifically highlight the differences in forecast accuracy attributable to the inclusion of luminosity function parameters, thereby emphasizing their significance in cosmological analyses. The fractional percentage differences (henceforth referred to as ‘differences’) for each parameter are presented in Fig. 7. The differences for the Year 1 (Y1) forecast are indicated by yellow squares, while those for Year 10 (Y10) are shown with pink circles. To aid in interpreting the plot, a line indicating a 20% difference is included, serving as a reference point to assess the significance of the differences observed.
@@
Figure 7: Fractional Differences between SRD and SRD+LF Scenarios Across Key Parameters. This figure illustrates the impact of including luminosity function parameters in intrinsic alignment (IA) modeling for an LSST cosmic shear analysis, quantified through fractional differences. Each symbol represents the fractional difference for a specific parameter: neon green squares for Year 1 (Y1) forecasts and neon teal circles for Year 10 (Y10) forecasts. The plot encompasses seven cosmological parameters and four IA parameters, showcasing their sensitivities to the inclusion of luminosity function parameters in different forecasting years. Notably, while some parameters like $\Omega_{b}$ and $h$ show negligible impact across both scenarios, significant effects are observed on $w_{0}$ and $w_{a}$ parameters, especially for the Y10 scenario. For the IA parameters, the fractional differences for $\beta$ and $\eta_{z}^{\mathrm{high}}$ highlight a substantial sensitivity to the inclusion of LF parameters, with $\beta$ and $\eta_{z}^{\mathrm{low}}$ showing a decrease in fractional differences. This decrease is interpreted as indicative of their heightened sensitivity to changes in luminosity function parameters, underscoring the critical role these parameters play in cosmological analyses for yielding accurate and comprehensive insights.
@@
This threshold, while arbitrary, provides a useful benchmark to gauge the impact of varying the luminosity function (LF) parameters in our analysis. Figure 7 displays the differences for seven cosmological parameters and four IA parameters across both forecasting years, serving as a quantitative measure of the change in parameter constraints. It’s important to note that LF parameters are not included in the plot, as the SRD (Mandelbaum et al., 2018) scenario does not vary these parameters, rendering a direct comparison impractical. However, in the SRD+LF scenario, these LF parameters are indeed varied, as reflected at the beginning of this section. Analysis of the plot reveals several key insights. As expected, the errors on the parameters $\Omega_{b}$ and $h$ are negligibly affected in both Y1 and Y10 forecasting scenarios. For parameters like $\Omega_{b}$, $\sigma_{8}$, and $n_{s}$, we observe that the errors remain modest for Y1. However, for Y10, where tighter constraints are expected, the fractional differences increase substantially. This observation is particularly significant for $w_{0}$ and $w_{a}$, parameters critical to the LSST’s science goals as a Stage IV cosmology survey. The sensitivity of these parameters to the expanded parameter space is notable. For Y10, the differences for dark energy equation of state parameters exceed 20%, impacting the survey’s objectives. Regarding the IA parameters, their sensitivity to the inclusion of LF parameters is evident, and expected considering their modeling involves the luminosity function. The parameter $\eta_{z}^{\mathrm{high}}$ shows differences below 10% across both forecasting years, aligning with our limited understanding of the IA amplitude at higher redshifts. For $A_{0}$ (a normalization factor), excluding LF parameters in the forecast appears to lead to about 16% overestimation of errors for Y1 and 19% Y10. The most pronounced impacts are observed in the scaling factor $\beta$ and $\eta_{z}^{\mathrm{high}}$. For the $\beta$ parameter, the error increases by 200% in both forecasting years. For $\eta_{z}^{\mathrm{high}}$, the increase is approximately 100% for Year 1 (Y1) and 80% for Year 10 (Y10). Notably, the degree of increase in uncertainty for these two parameters shows a relative improvement in Year 10 compared to Year 1. This could be potentially attributed to the increased volume of data available in the later observing stages of the LSST, which offers more overall constraining power.
@@
This preliminary investigation underscores the crucial role and potential of fully integrating luminosity function parameters into cosmological analyses. Jointly modeling systematic effects via a shared luminosity function not only leverages underlying correlations but also enriches our understanding of the interconnections among various parameters. For completeness, we include the Fisher matrix marginalized contours in the Appendix in Fig. 11.
4.4 Forecasts for Luminosity-Function-Based Joint Modeling of IA and N(z)
@@
Having established the qualitative effect of varying the luminosity function parameters on the convolutional modeling method for redshift distributions (Sec. 4.1), the data vector (Sec. 4.2), and explored the impact on forecast constraints of varying the LF parameters in the IA modeling only (Sec. 4.3), we now address the primary question of this work: what is the impact of jointly modeling IA and the redshift distribution via the luminosity function on cosmological parameter constraints?
@@
In this study, we analyse four distinct Fisher forecasting scenarios: two correspond to our integrated framework for modeling astrophysical systematics in the LSST Year 1 and Year 10 forecasts (denoted as JMAS Y1 and JMAS Y10, respectively), and two are based on a standard analysis approach, which treats intrinsic alignment (IA) and redshift distributions independently, as outlined in Mandelbaum et al. (2018), for LSST Year 1 and Year 10 (referred to as SRD Y1 and SRD Y10, respectively). The forecasted posterior parameter constraints on cosmological parameters for LSST Year 1 and LSST Year 10 are shown in Fig. 8 (the complete parameter space, encompassing the IA and LF parameters, will be thoroughly explored in subsequent sections). LSST Year 1 posterior constraints are illustrated in the lower segment of the triangle plot in Fig. 8, while LSST Year 10 constraints are presented in the upper corner of the figure. The JMAS Y1 constraints are visualized using light green filled contours, while the SRD Y1 constraints are shown as unfilled dashed contours in dark green. For Year 10, the JMAS Y10 scenario is represented with blue filled contours, and the SRD Y10 scenario with unfilled dashed contours in dark blue. The diagonal elements present 1D marginal distributions on parameters across all scenarios.
@@
Our key objective is to assess the effect on the constraints of cosmological parameters from 1) decoupling the volume selection from the luminosity function selection in the redshift distribution modeling and 2) considering the latter jointly with intrinsic alignment, especially when the parameter space is expanded. This analysis is crucial, as subtle differences in modeling IA and $N(z)$, even with negligible variations in the redshift distribution, can significantly influence the inferred cosmological parameters. The sensitivity of cosmic shear to these modeling details is highlighted by the notable variance in constraining power and the change in degeneracy directions. This is particularly apparent in the flipped correlation between $n_{s}$ and both $\Omega_{m}$ and $\sigma_{8}$ parameters, a reversal distinctly showcased in Fig. 8. Such inversions in degeneracy directions can be attributed to the influence of newly-introduced nuisance luminosity function parameters, which interact with cosmological parameters.
@@
Despite this increased variance in constraining power, our approach achieves cosmological parameter constraints which are comparable with, if somewhat broader than, cosmological forecasts using more standard methods. These broader constraints can be viewed to some extent as a more realistic representation of the inherent uncertainties, particularly in relation to the use of a luminosity-function-dependent IA model in both cases. It is in this context of that we delve now into the detailed examination of parameter degeneracies and potential biases, providing a comprehensive understanding of the implications of our modeling choices.
@@
Figure 8: Fisher matrix marginalized contours for forecasts across different analysis scenarios. This figure depicts relationships between JMAS (this work) Y1 and SRD Y1 (lower triangle) and JMAS Y10 ad SRD Y10 (upper triangle). The plots contrast the effects of decoupling volume selection and luminosity function on cosmological parameters, with green filled contours representing JMAS Y1, unfilled dark green contours for SRD Y1 (dashed), blue filled contours for JMAS Y10, and unfilled blue dashed contours for SRD Y10. The diagonal elements showcase 1D marginal distributions across all scenarios. No significant variations in constraining power and degeneracy directions are present, emphasizing the ability of our joint modelling method to successfully reproduce the forecasting scenarios.
@@ 4.4.1 Degeneracies: cosmological and luminosity function parameters
Expanding on our earlier discussions, this section explores the degeneracies observed between some key cosmological parameters and the parameters of the luminosity function. The comprehensive corner plots for the complete parameter space, pertaining to the forecast years 1 and 10, are presented in the Appendix, specifically in Figures 13 and 14. For a focused analysis, we have extracted and highlighted the segments showcasing the interplay between the cosmological and luminosity function parameters. These specific segments are presented for comparison in Figure 9, offering a clearer view of their relationships.
@@
Spectral index $\boldsymbol{n_{s}}$. In our Fisher analysis with the luminosity function joint modeling approach, we observe consistent but wider constraints between the spectral index $n_{s}$ and other cosmological parameters. Notably, there is a reversal in the directions of the ellipses for marginalised 2D constraints of $n_{s}$ and both $\Omega_{m}$ and $\sigma_{8}$. This reversal is likely influenced by the specific modeling choices and fiducial values of the luminosity function (LF) parameters. While $n_{s}$ remains generally poorly constrained by cosmic shear data, this observation provides valuable insights into the impact of LF modeling on the analysis, highlighting the intricate interplay between LF modeling and cosmological parameter estimation.
@@
The evolution in some cases of these correlations between Y1 and Y10 highlights the necessity of continually updating our models and assumptions with the progression of observational data. We note that is is important to bear in mind, when interpreting these results, that fiducial luminosity function parameters differ between Y1 and Y10. This distinction implies luminosity function parameter constraints will showcase disparate degeneracies and magnitudes, given they aren’t identical random variables. That is to say: we do not expect the values of the true luminosity function parameters for LSST Y1 to be the same as for Y10.
@@ 4.4.2 Degeneracies: intrinsic alignment and luminosity function parameters
We now examine the interplay and degeneracies between intrinsic alignment (IA) parameters and the luminosity function parameters.
Overall intrinsic alignment amplitude $\boldsymbol{A_{0}}$. The intrinsic alignment (IA) signal is generally stronger in denser environments, where gravitational interactions and tidal forces that contribute to intrinsic alignments are more pronounced. However, this trend does not uniformly apply to all galaxy types. For example, satellite galaxies often exhibit minimal alignment. Studies by Sifón et al. 2015, Chisari et al. 2015, and Georgiou et al. 2019 indicate that while alignments in higher density areas tend to be stronger, this pattern does not necessarily hold for satellite galaxies, pointing to a complex interplay of factors. The observed anti-correlation between $A_{0}$ and $\phi_{0}^{*}$, where $A_{0}$ serves as a scaling factor adjusting the IA signal strength and $\phi_{0}^{*}$ reflects the density of source galaxies at a given redshift, illustrates this complexity. The direct multiplication of $A_{0}$ and $\phi_{0}^{*}$ within the IA model used explains their negative correlation, offering a straightforward insight into the dynamics of intrinsic alignments in relation to galaxy density. Higher values of $\phi_{0}^{*}$ suggest a greater density of source galaxies at a particular redshift, which could logically lead to enhanced clustering. This in turn is likely to result in a stronger intrinsic alignment signal. Thus, increasing $\phi_{0}^{*}$ can be compensated by decreasing $A_{0}$.
@@
Power-law luminosity scaling $\boldsymbol{\beta}$. As $\beta$ is the power law index of the luminosity-dependent term in the IA amplitude, we would expect it to be degenerate with the parameters of our luminosity function. In Year 1, we see this to be generically true. However, all but the correlation with $\phi_{0}^{*}$ effectively vanish in Year 10, which is quite surprising. This might indicate a complex interplay between galaxy properties and the intrinsic alignments that evolves with the statistical power of our data, and moreover is potentially influenced by the specific adjustments made to the fiducial LF parameters to match the Y1 and Y10 redshift distribution as appropriate. This discrepancy between Y1 and Y10 highlights the necessity of considering long-term observational data.

### v3

@@ 4.3 Impact of varying luminosity function parameters in IA modelling only
Before considering a full joint modelling of IA and source photometric redshift distributions via the luminosity function, we consider also the impact of simply marginalizing over the luminosity function parameters of the IA modelling an LSST cosmic shear analysis. In the case of the LSST Science Requirements Document (SRD) analysis of Mandelbaum et al. (2018) (which we use extensively in this work as a comparison point and an example of a standard analysis set-up), a Schechter luminosity function with fixed parameter values from the GAMA and DEEP2 surveys is used. The idea of including luminosity function parameters as nuisance IA modelling parameters was explored in Krause et al. (2016) for an LSST-like survey; here we present the results of doing so in our particular set-up to build intuition prior to introducing the luminosity function jointly within the redshift distribution and IA modelling.
@@
Since we aim to understand the importance of including luminosity function parameters in the forecast, we analyze the resulting Fisher matrices to extract the $1-\sigma$ errors on parameters. To quantitatively assess the impact of this inclusion, we compute the fractional difference between the SRD and SRD+LF scenarios for these errors on the parameters. This is defined as:
@@
In Eq. 30, $\sigma$ represents a $1\sigma$ error on parameter $\theta$, where $X$ and $Y$ denote the forecasting scenarios, with $Y$ being the benchmark. This approach allows us to specifically highlight the differences in forecast accuracy attributable to the inclusion of luminosity function parameters. The fractional percentage differences (henceforth referred to as ‘differences’) for each parameter are presented in Fig. 7. The differences for the Year 1 (Y1) forecast are indicated by green squares, while those for Year 10 (Y10) are shown with teal circles. To aid in interpreting the plot, a line indicating a 20% difference is included.
@@
Figure 7: Fractional Differences between SRD and SRD+LF Scenarios Across Key Parameters. This figure illustrates the impact of including luminosity function parameters in intrinsic alignment (IA) modeling for an LSST cosmic shear analysis, quantified through fractional differences in marginalized $1\sigma$ errors. Each symbol represents the fractional difference for a specific parameter: neon green squares for Year 1 (Y1) forecasts and neon teal circles for Year 10 (Y10) forecasts. The plot encompasses seven cosmological parameters and four IA parameters, showcasing their sensitivities to the inclusion of luminosity function parameters in different forecasting years. Notably, while some parameters like $\Omega_{b}$ and $h$ show negligible impact across both scenarios, significant effects are observed on $w_{0}$ and $w_{a}$ parameters, especially for the Y10 scenario. For the IA parameters, the fractional differences for $\beta$ and $\eta_{z}^{\mathrm{high}}$ highlight a substantial sensitivity to the inclusion of LF parameters, with $\beta$ and $\eta_{z}^{\mathrm{low}}$ showing a decrease in fractional differences between Y1 and Y10.
@@
Figure 7 displays the differences for seven cosmological parameters and four IA parameters across both forecasting years. It is important to note that LF parameters are not included in the plot, as the SRD (Mandelbaum et al., 2018) scenario does not vary these parameters. However, in the SRD+LF scenario, these LF parameters are indeed varied, as reflected at the beginning of this section. Analysis of the plot reveals several key insights. As expected, the errors on the parameters $\Omega_{b}$ and $h$ are negligibly affected in both Y1 and Y10 forecasting scenarios. For parameters like $\Omega_{b}$, $\sigma_{8}$, and $n_{s}$, we observe that the errors remain modest for Y1. However, for Y10, where tighter constraints are expected, the fractional differences increase substantially. This observation is particularly significant for $w_{0}$ and $w_{a}$, parameters critical to the LSST’s science goals as a Stage IV cosmology survey. The sensitivity of these parameters to the expanded parameter space is notable. For Y10, the differences for dark energy equation of state parameters exceed 20%, impacting the survey’s objectives. Regarding the IA parameters, their sensitivity to the inclusion of LF parameters is evident, and expected considering their modelling involves the luminosity function. The parameter $\eta_{z}^{\mathrm{high}}$ shows differences below 10% across both forecasting years, aligning with our limited understanding of the IA amplitude at higher redshifts. For $A_{0}$, including LF parameters in the forecast appears to lead to about 16% increase in errors for Y1 and 19% Y10. The most pronounced impacts are observed in the scaling factor $\beta$ and $\eta_{z}^{\mathrm{high}}$. For the $\beta$ parameter, the error increases by 200% in both forecasting years. For $\eta_{z}^{\mathrm{high}}$, the increase is approximately 100% for Year 1 (Y1) and 80% for Year 10 (Y10). Notably, the degree of increase in uncertainty for these two parameters shows a relative improvement in Year 10 compared to Year 1. This could be potentially attributed to the increased volume of data available in the later observing stages of the LSST, which offers more overall constraining power.
@@
This preliminary investigation underscores the potential benefit of fully integrating luminosity function parameters into cosmological analyses. Jointly modelling systematic effects via a shared luminosity function not only leverages underlying correlations but also enriches our understanding of the interconnections among various parameters. For completeness, we include the Fisher matrix marginalized contours in the Appendix in Fig. 13.
4.4 Forecasts for Luminosity-Function-Based Joint Modelling of IA and N(z)
@@
Having established the qualitative effect of varying the luminosity function parameters on the modellng method for redshift distributions (Sec. 4.1), the data vector (Sec. 4.2), and explored the impact on forecast constraints of varying the LF parameters in the IA modelling only (Sec. 4.3), we now address the primary question of this work: what is the impact of jointly modelling IA and the redshift distribution via the luminosity function on cosmological parameter constraints?
@@
In this study, we analyze four distinct Fisher forecasting scenarios: two correspond to our integrated framework for modelling astrophysical systematics in the LSST Year 1 and Year 10 forecasts (denoted as JMAS Y1 and JMAS Y10, respectively), and two are based on a standard analysis approach, which treats intrinsic alignment (IA) and redshift distributions independently, as outlined in Mandelbaum et al. (2018), for LSST Year 1 and Year 10 (referred to as SRD Y1 and SRD Y10, respectively). The forecasted parameter constraints on cosmological parameters for LSST Year 1 and LSST Year 10 are shown in Fig. 8 (the complete parameter space, encompassing the IA and LF parameters, will be thoroughly explored in subsequent sections). LSST Year 1 posterior constraints are illustrated in the lower segment of the triangle plot in Fig. 8, while LSST Year 10 constraints are presented in the upper corner of the figure. The JMAS Y1 constraints are visualized using light green filled contours, while the SRD Y1 constraints are shown as unfilled dashed contours in dark green. For Year 10, the JMAS Y10 scenario is represented with blue filled contours, and the SRD Y10 scenario with unfilled dashed contours in dark blue. The diagonal elements present 1D marginal distributions on parameters across all scenarios.
@@
Our key objective is to assess the effect on the constraints of cosmological parameters from 1) decoupling the volume selection from the luminosity function selection in the redshift distribution modeling and 2) considering the latter jointly with intrinsic alignment, especially when the parameter space is expanded. The sensitivity of cosmic shear to these modelling details is highlighted by the notable variance in constraining power and the change in degeneracy directions between the JMAS and SRD cases. This is particularly apparent in the flipped correlation between $n_{s}$ and both $\Omega_{m}$ and $\sigma_{8}$ parameters see Fig. 8. Such inversions in degeneracy directions can be attributed to the influence of newly introduced nuisance luminosity function parameters, which interact with cosmological parameters.
@@
Despite this, our approach achieves cosmological parameter constraints which are comparable with, if somewhat broader than, cosmological forecasts using more standard methods. These broader constraints can be viewed to some extent as a more realistic representation of the inherent uncertainties, particularly in relation to the use of a luminosity-function-dependent IA model in both cases. It is in this context that we delve now into the detailed examination of parameter degeneracies and potential biases.
@@
Figure 8: Fisher matrix marginalized contours for forecasts across different analysis scenarios. This figure depicts relationships between JMAS (this work) Y1 and SRD Y1 (lower triangle) and JMAS Y10 ad SRD Y10 (upper triangle). The plots contrast the effects of decoupling volume selection and luminosity function on cosmological parameters, with green filled contours representing JMAS Y1, unfilled dark green contours for SRD Y1 (dashed), blue filled contours for JMAS Y10, and unfilled blue dashed contours for SRD Y10. The diagonal elements showcase 1D marginal distributions across all scenarios.
@@ 4.4.1 Degeneracies: cosmological and luminosity function parameters
This section explores the degeneracies observed between some key cosmological parameters and the parameters of the luminosity function. The corner plots for the complete parameter space, pertaining to the forecast years 1 and 10, are presented in the Appendix, specifically in Figures 14 and 15. For a focused analysis, we have extracted and highlighted the segments showcasing the interplay between the cosmological and luminosity function parameters. These specific segments are presented for comparison in Figure 9.
@@
Spectral index $\boldsymbol{n_{s}}$. In our Fisher analysis with the luminosity function joint modelling approach, we observe consistent but wider constraints between the spectral index $n_{s}$ and other cosmological parameters. Notably, there is a reversal in the directions of the ellipses for marginalised 2D constraints of $n_{s}$ and both $\Omega_{m}$ and $\sigma_{8}$. This reversal is likely influenced by the specific modelling choices and fiducial values of the luminosity function (LF) parameters. While $n_{s}$ remains generally poorly constrained by cosmic shear data, this observation provides valuable insights into the impact of LF modelling on the analysis, highlighting the intricate interplay between LF modelling and cosmological parameter estimation.
@@
The evolution in some cases of these correlations between Y1 and Y10 highlights the necessity of continually updating our models and assumptions with the progression of observational data. We note that it is important to bear in mind, when interpreting these results, that fiducial luminosity function parameters differ between Y1 and Y10. This distinction implies luminosity function parameter constraints will showcase disparate degeneracies and magnitudes, given they are not identical random variables. That is to say: we do not expect the values of the true luminosity function parameters for LSST Y1 to be the same as for Y10.
@@ 4.4.2 Degeneracies: intrinsic alignment and luminosity function parameters
We now examine the interplay and degeneracies between intrinsic alignment (IA) parameters and the luminosity function parameters. These correlations are given in Figure 10.
Overall intrinsic alignment amplitude $\boldsymbol{A_{0}}$. The observed anti-correlation between $A_{0}$ and $\phi_{0}^{*}$ illustrates the direct relationship between the IA signal strength and the fractional number density of galaxies subject to IA effects (in this work: red galaxies). The direct multiplication of $A_{0}$ and $\phi_{0}^{*}$ within the IA model (see Eq. 9) explains their negative correlation: increasing $\phi_{0}^{*}$, and hence the fraction of red galaxies, can be compensated by decreasing the overall IA strength $A_{0}$.
@@
(a) Cosmological and LF parameters for LSST Y1.
(b) Cosmological and LF parameters for LSST Y10.
Figure 9: Segments of the Fisher matrix marginalized contours for JMAS analysis across forecasting years. The left panel illustrates the correlations between cosmological and LF parameters for LSST Year 1 (Y1, green contours), while the right panel shows the same for Year 10 (Y10, blue contours). This side-by-side arrangement highlights the progression of parameter constraints over time.
(a) IA and LF parameters for LSST Y1.
(b) IA and LF parameters for LSST Y10.
Figure 10: Segments of the Fisher matrix marginalized contours for JMAS analysis across forecasting years. The left panel illustrates the correlations between IA and LF parameters for LSST Year 1 (Y1, green contours), while the right panel shows the same for Year 10 (Y10, blue contours). This side-by-side arrangement highlights the progression of parameter constraints over time.
(a) LF parameters for LSST Y1.
(b) LF parameters for LSST Y10.
Figure 11: Segments of the Fisher matrix marginalized contours for JMAS analysis across forecasting years. The left panel illustrates the correlations between the LF parameters for LSST Year 1 (Y1, green contours), while the right panel shows the same for Year 10 (Y10, blue contours). This side-by-side arrangement highlights the progression of parameter constraints over time.
Power-law luminosity scaling $\boldsymbol{\beta}$. As $\beta$ is the power law index of the luminosity-dependent term in the IA amplitude, we would expect it to be degenerate with the parameters of our luminosity function. In Year 1, we see this to be generically true. However, all but the correlation with $\phi_{0}^{*}$ effectively vanish in Year 10, which is quite surprising. This might indicate a complex interplay between galaxy properties and the intrinsic alignments that evolves with the statistical power of our data, and moreover is potentially influenced by the specific adjustments made to the fiducial LF parameters to match the Y1 and Y10 redshift distribution as appropriate.